if measure of angle A equals 3x+42, measure of angle B=6x+18, angle A and angle C are supplementary, and angle B and angle C are supplementary find x

Answer by ewatrrr(24785)

Hi,
.
Given:
angle A and angle C are supplementary,
and angle B and angle C are supplementary
.
Definition of supplements
A + C = 180
B + C = 180
.
Transitive Property of Equality
A + C = B + C
.
Subtraction Property of Equality
A + C - C = B + C - C
.
Remaining from Above
Conclusion:
A = B
.
A= 3x+42 and B=6x+18
.
3x + 42 = 6x + 18
.
simpliy and solve
24 = 3x
8 = x
.
x = 8 degrees
.
check your answer
24 + 42 = 48 + 18 = 66